OpenStudy (anonymous):

How will the equation y= -x2 + 3 compare to the original parent function (y=x2)?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

im thinking it could be slides up

OpenStudy (skullpatrol):

$$\Huge y=-x^2+3$$ $$\Huge (y=x^2)$$

OpenStudy (skullpatrol):

plot the points -1, 0, 1 for each

OpenStudy (anonymous):

i know already did im just wondering if my answers right

OpenStudy (skullpatrol):

what is your answer?

OpenStudy (anonymous):

shift/slides up

OpenStudy (skullpatrol):

what effect does the lowered minus sign (-) have?

OpenStudy (anonymous):

?

OpenStudy (skullpatrol):

compare y = x^2 to y = -x^2

OpenStudy (zale101):

 The graph, y=x^2 is a parabola :)

OpenStudy (skullpatrol):

note y = -x^2 means y = -(x^2)

OpenStudy (zale101):

there's a big change in the graph once the whole equation is multiplied by -1